OpenStudy (anonymous):

integrate tan^4(x)sec(x)dx

OpenStudy (anonymous):

I have it to a point where it's going to get really long...I think I'm having trouble with the reduction formulas (I have it split into 3 integrals right now)

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\sec^5xdx -2\int\limits_{?}^{?}\sec^3xdx + \int\limits_{?}^{?}secxdx\] nevermind the ?'s on the intergrals...not sure how to get rid of them

OpenStudy (jamesj):

I would write tan^4 = (sec^2 - 1)^2 Then expand the integral out into a number of integrals involving sec^n Then find each of these in turn. You may find it helpful to use (and derive if needs be) the standard reduction for formula for \[ I_n = \int \sec^n x \ dx = stuff + coefficient.\int \sec^{n-2} x \ dx \]

OpenStudy (anonymous):

@JamesJ I wonder why don't we let u = tanx, ---> du = tanx secx dx and we break the original function to \[\int\limits_{}^{} \tan^3 tanxsecxdx\] and then replace as \[\int\limits_{}^{} u^3 du = 3u^2 +C \] and replace back to x to get \[3\tan^3x + C\].is it wrong?

OpenStudy (jamesj):

yes, it's wrong. The derivative of tan x is sec^2 x

OpenStudy (anonymous):

thanks i got it :)

OpenStudy (anonymous):

Thanks guys, forgot I asked this. Went back and looked at it and used reduction formulas (2 different ones for the sec^5 integral and the sec^3 integral) and just used the integral on the last (sec(x))...took a good sheet of paper but after combining like terms after I got the answers from the formulas, I came out with the right answer :D